OpenStudy (samigupta8):

There are four capacitors with capacitances and breakdown voltages mentioned. Two of them are in series (3C,1kV and 2C,2kV) and this is in parallel with two other capacitor in series (7C,1kV and 3C,2kV) .What should be the maximum value of external emf source such that no capacitor breaks down??

OpenStudy (samigupta8):

@ganeshie8 pls help

OpenStudy (samigupta8):

@priyar

OpenStudy (samigupta8):

@ganeshie8

ganeshie8 (ganeshie8):

Q is same on capacitors are in series

ganeshie8 (ganeshie8):

Say the voltages across capacitors in given order are V1, V2, V3, V4

ganeshie8 (ganeshie8):

Let V be the maximum emf allowed

ganeshie8 (ganeshie8):

Then you need to find V such that : V = V1 + V2 V = V3 + V4 3CV1 = 2CV2 7CV3 = 3CV4 V1 < 1kV V2 < 2kV V3 < 1kV V4 < 2kV

OpenStudy (samigupta8):

Correct !!

OpenStudy (samigupta8):

Does that mean that V=5/3V2 And V=10/3V3??

OpenStudy (samigupta8):

So potential should be 10/3 .. Ryt??

OpenStudy (samigupta8):

Bt that's not the ans!

ganeshie8 (ganeshie8):

express each of V1, V2, V3, V4 in terms of V

OpenStudy (samigupta8):

Or even the opposite of it.. V in terms of each of them..!

OpenStudy (samigupta8):

And solving this i got V can be at max 10/3..!

ganeshie8 (ganeshie8):

is the answer 2.5kV ?

OpenStudy (samigupta8):

Lemme check!

OpenStudy (samigupta8):

Yep..

ganeshie8 (ganeshie8):

then try again

ganeshie8 (ganeshie8):

i have got that answer using your method oly

OpenStudy (samigupta8):

Okay..

OpenStudy (samigupta8):

I must see once again then!

ganeshie8 (ganeshie8):

V < 5/3V2 V < 2/5V1 V < 10/3V3 V < 10/7V4

OpenStudy (samigupta8):

Ty!!

ganeshie8 (ganeshie8):

Np:)